OpenStudy (cherrilyn):

Determine convergence or divergence and explain which method you used for......

OpenStudy (cherrilyn):

\[\sum_{n=2}^{\infty} 1/ n \ln n-n\]

OpenStudy (anonymous):

the whole term in in denominator?

OpenStudy (cherrilyn):

no.. sorry I don't know how to make a division sign on here...... but its 1 over nln-n

OpenStudy (anonymous):

thats what i want to confirm

OpenStudy (cherrilyn):

yeah. 1 is in the numerator.

OpenStudy (anonymous):

have u read the integral test for convergence of the series?

OpenStudy (cherrilyn):

If \[\int\limits_{1}^{\infty} f(x)dx\] converges then \[\sum_{n=1}^{\infty} a _{n} \] converges

OpenStudy (anonymous):

right and function must be montonically decreasing

OpenStudy (anonymous):

\[\int\limits_{2}^{\infty} (1/x)/(\ln x-1) dx\]

OpenStudy (cherrilyn):

?

OpenStudy (anonymous):

solve the integral, the given series converges if the integral converges

OpenStudy (anonymous):

\[\lim t \rightarrow \infty [\ln(lnx-1)] x varies from 2 \to t\]